OpenStudy (anonymous):

*I will give you meddles and become you're fan, just can someone please help me with this question? I've been stuck on this for like a hour now.* Use the chart below to Find the equation of the line in the graph, and mathematically model the scenario using function notation where cost is a function of number of minutes.

OpenStudy (anonymous):

@ganeshie8 @geerky42 @Hero

jimthompson5910 (jim_thompson5910):

can you identify two points on that blue line?

OpenStudy (anonymous):

(0,20) & (700,50) ?

jimthompson5910 (jim_thompson5910):

I agree with (0,20) but (700,50) is a bit harder to agree with

jimthompson5910 (jim_thompson5910):

for me, I would say (0,20) and (400,40)

OpenStudy (anonymous):

Okay, now i always get stuck while trying to put it into a equation.

jimthompson5910 (jim_thompson5910):

after you get these two points, you find the slope m = (y2 - y1)/(x2 - x1) m = (40 - 20)/(400-0) m = 20/400 m = 1/20 Hopefully you can see how I get the slope of 1/20 ?

OpenStudy (anonymous):

when its at 20/40 did you divide by 20?

jimthompson5910 (jim_thompson5910):

20/400 reduces to 1/20 you divide both parts by the GCF 20

OpenStudy (anonymous):

alright, now that we have 1/20 what do we do?

jimthompson5910 (jim_thompson5910):

then we use this slope with either point to get the y-intercept let's say we use the point (400,40) y = mx + b y = (1/20)*x + b 40 = (1/20)*(400) + b 40 = 20 + b 40 - 20 = b 20 = b b = 20 So the y-intercept is 20 And you can look at the point (0,20) that is on the line as well. This point immediately tells you what the y-intercept is because x = 0 here and it crosses the y axis at this point

jimthompson5910 (jim_thompson5910):

So to recap, we got m = 1/20 (slope) b = 20 (y-intercept) So... \[\Large y = mx + b\] \[\Large y = \frac{1}{20}x + b\] \[\Large y = \frac{1}{20}x + 20\] is your equation

jimthompson5910 (jim_thompson5910):

hopefully it all makes sense?

OpenStudy (anonymous):

The equation I got; y = 1/20x + 20 where x is minutes and f(x) is the cost in dollars.

jimthompson5910 (jim_thompson5910):

that is correct, the answer is [\Large y = \frac{1}{20}x + 20\] or [\Large f(x) = \frac{1}{20}x + 20\]

jimthompson5910 (jim_thompson5910):

x is the number of minutes y or f(x) is the cost in dollars

OpenStudy (anonymous):

Thank you so much, you helped me so much!!(:

OpenStudy (anonymous):

Can you maybe help me on another one? (:

jimthompson5910 (jim_thompson5910):

I'm glad I have and sure, I can do another